# LAG: LLM agents for Leaderboard Auto Generation on Demanding

### v1

@@ 3.3 Table Unpacking and Integration
Following the table extraction and classification phase, each table $Table_{i}$ is sent into the LLM to extract the core information. To build a useful and high-quality leaderboard, we define four types of scientific terms: Datasets, Metrics, Experiment Results, and Experiment Settings. For datasets, we use LLMs to count the frequency in all filtered papers $P_{\text{filtered}}$ of each dataset under a certain research topic and retain the top-K (K=5) datasets with the highest frequency of occurrence in scientific papers. For the rest of the three scientific terms, we utilize LLMs to extract from given $Table_{i}$ with a related table description $D_{i}$. After scientific term extraction, we recombined them into a quintuple, including the paper title as the unique identification ID. Each paper can produce one quintuple and finally we get a raw leaderboard with $M$ quintuples from $M$ filtered papers. The raw leaderboard is reranked on the basis of the experiment results.
3.4 Leaderboard Generation and Evaluation
@@
After we obtain $K$ leaderboards based on top-K frequent datasets, the final stage involves a quality evaluation based on our pre-defined four criteria, which is shown in Table 4 in Appendix. Each leaderboard is assigned three scores based on “Coverage”, “Latest” and “Structure”. Since a research topic may contain several datasets, the “Multi-Aspect” is the average quality score that is used to evaluate the LLM-generated leaderboards for each dataset. The best leaderboard is chosen from $N$ candidates. LLMs critically examine the leaderboards in several aspects. The final output of Leaderboard is $L_{\text{best}}=\text{Evaluate}({L_{\text{ca1}},L_{\text{ca2}},...,L_{\text{caN}}})$.
The methodology outlined here, from paper collection to leaderboard evaluation, ensures that LAG effectively addresses the complexities of leaderboard generation in the AI domain using advanced LLM agents. We provide Pseudo-code for easily understanding, which is shown in Algorithm 1.
@@
1:  Input: Scientific topic $T$, open-access platform arXiv $A$
2:  Output: Final refined and evaluated leaderboard $L$ Stage 1: Paper Collection and Document Split
3:  Crawl topic $T$ related $N$ publications $P_{\text{init}}=\{P_{1},...P_{N}\}\leftarrow\text{Retrieve}(T,A)$
4:  Filter out topic-unrelated and old papers, $P_{\text{filtered}}=\{P_{1},...P_{M}\}\leftarrow\text{Retrieve}(P_{\text{init}},\text{date},\text{topic})$ Stage 2: Table Extraction and Classification
5:  for each Leaderboard iteration $i=1$ to $Iters$ do
@@
13:        Recombine all quintuples and rank the quintuples by performance scores.
@@
Leaderboard Length (items) Topic-related Quality Model Speed/s Content Quality Recall Precision Coverage Latest Structure Multiaspect 5 $76.57_{\pm 11.65}$ $79.43_{\pm 8.86}$ Qwen2.5-7B $131.43$ $3.60_{\pm 0.48}$ $3.46_{\pm 0.49}$ $3.18_{\pm 0.32}$ $3.41$ Qwen2.5-14B $129.51$ $4.23_{\pm 0.38}$ $4.14_{\pm 0.31}$ $3.68_{\pm 0.29}$ $4.02$ GPT4-o $49.64$ $4.52_{\pm 0.42}$ $4.70_{\pm 0.32}$ $4.32_{\pm 0.38}$ $4.41$ O1-preview $79.67$ $4.63_{\pm 0.48}$ $4.71_{\pm 0.71}$ $4.40_{\pm 0.33}$ $4.58$ Human Writing $355$ 4.89 4.83 4.91 4.88 10 $75.19_{\pm 9.81}$ $80.05_{\pm 6.76}$ Qwen2.5-7B $156.41$ $3.22_{\pm 0.48}$ $3.41_{\pm 0.49}$ $4.11_{\pm 0.39}$ $3.57$ Qwen2.5-14B $163.54$ $3.91_{\pm 0.48}$ $3.55_{\pm 0.49}$ $3.41_{\pm 0.39}$ $4.61$ GPT4-o $88.96$ $4.68_{\pm 0.39}$ $4.59_{\pm 0.33}$ $4.45_{\pm 0.41}$ $4.56$ O1-preview $98.44$ $4.40_{\pm 0.48}$ $4.46_{\pm 0.71}$ $4.31_{\pm 0.33}$ $4.39$ Human Writing $612$ $4.81$ $4.72$ $4.65$ $4.72$ 15 $71.34_{\pm 8.39}$ $74.58_{\pm 7.35}$ Qwen2.5-7B $183.45$ $3.11_{\pm 0.28}$ $3.23_{\pm 0.26}$ $3.15_{\pm 0.27}$ $3.16$ Qwen2.5-14B $195.63$ $3.68_{\pm 0.28}$ $3.32_{\pm 0.19}$ $3.18_{\pm 0.24}$ $3.39$ GPT4-o $105.61$ $4.47_{\pm 0.22}$ $4.16_{\pm 0.27}$ $4.32_{\pm 0.24}$ $4.28$ O1-preview $109.33$ $4.21_{\pm 0.48}$ $4.06_{\pm 0.21}$ $4.28_{\pm 0.31}$ $4.18$ Human Writing $839$ $4.71$ $4.65$ $4.44$ $4.60$ 20 $67.58_{\pm 9.12}$ $70.33_{\pm 6.89}$ Qwen2.5-7B $196.33$ $3.03_{\pm 0.25}$ $3.11_{\pm 0.31}$ $2.98_{\pm 0.25}$ $3.16$ Qwen2.5-14B $208.64$ $3.49_{\pm 0.34}$ $3.17_{\pm 0.26}$ $3.03_{\pm 0.28}$ $3.39$ GPT4-o $120.52$ $4.28_{\pm 0.28}$ $3.92_{\pm 0.22}$ $4.21_{\pm 0.25}$ $4.13$ O1-preview $117.45$ $4.12_{\pm 0.38}$ $3.96_{\pm 0.25}$ $4.16_{\pm 0.29}$ $4.08$ Human Writing $1128$ $4.72$ $4.68$ $4.34$ $4.58$
Table 2: Results of leaderboard quality generated by LLMs in the first iteration. Leaderboard Length: The number of items in the leaderboard. For example, a 5-item leaderboard contains 5 baselines. Topic-related Quality: The precision and recall of each paper in relation to its relevance to the topic. Speed: The average time required to generate a single leaderboard. Content Quality: The evaluation results of the leaderboard content.
4 Experiments
We designed experiments for LAG, aiming to answer four questions: RQ-1: Can LAG address the paper coverage issue and generate fair leaderboards by incorporating the latest baselines? RQ-2: Can LAG reduce time consumption? RQ-3: Is the evaluation consistent between LAG and human experts? RQ-4: Is each proposed component of LAG useful?
4.1 Experimental Setup
We evaluated LAG’s performance by testing its ability to generate leaderboards for specific topics across various complex settings.
4.1.1 Evaluation Metrics
We use two metrics to evaluate the quality (topic-related and leaderboard content) and speed of leaderboard generation, respectively, in response to the three challenges mentioned in the introduction.
(1) Topic-related Quality: The aforementioned arXiv crawler employs regular expression matching in the abstract section to identify papers related to specified topics. While this method is efficient, it is relatively rudimentary and cannot guarantee that all retrieved papers meet our requirements. The quality of these papers not only directly affects the final leaderboard, but low-quality candidate papers can also significantly prolong the time required for construction. Therefore, it is essential to evaluate the quality of the retrieved articles. We evaluate the quality of content from the following two aspects. (i) Recall: It measures whether all items in the generated leaderboard are related to the given research topic. (ii) Precision: It identifies irrelevant items, ensuring that the items in the generated leaderboards are pertinent and directly support the given research topic.
(2) Leaderboard Content Quality: The evaluation metric of leaderboard content quality includes four aspects. Each aspect is judged by LLMs according to a 5-point, calibrated by human experts. The evaluation criteria are listed in Table 4. (i) Coverage: Assess each paper represented on the LAG-generated leaderboards encapsulates all aspects of the topic. (ii) Latest: Test whether all papers represented on the LAG-generated leaderboards are latest. (iii) Structure: Evaluate the logical organization and determine whether LAG leaderboards are missing any items. (iv) Multiaspect: Average score of the previous three criteria for LAG-generated leaderboards.
(3) Leaderboard Construction Speed: Manually building a leaderboard is a time-consuming and laborious task. This process can be divided into the following main components: $T_{r}$ (search for papers on a specific topic), $T_{b}$ (browse all retrieved articles and develop several highly frequent datasets), $T_{f}$ (filter candidate articles based on the selected datasets), $T_{e}$ (read and extract information), and $T_{c}$ (the integration and construction time). And the total time consumption can be calculated as:
$T_{\text{manual}}=T_{r}+T_{b}+T_{f}+T_{e}+T_{c}.$ (1)
Given $L$ denotes the length of the leaderboard, $N_{\text{retrieved}}$ number of retrieved articles, $N_{\text{filtered}}$ number of articles retained, and $P$ the proportion of valid articles with $P=\frac{N_{\text{filtered}}}{N_{\text{retrieved}}}$. We find that $T_{b}$ and $T_{f}$ are strongly correlated with leaderboard length $L$ and the Topic-related quality:
While $T_{r}$ is relatively fixed, $T_{e}$ and $T_{c}$ usually only have a positive correlation with $L$.
For LAG, we barely account for all the invocation time of the agents’ API calls. Compared to manual work, which often takes several days, LAG reduces the total time cost in the minute level. This is largely attributed to the task decomposition conducted in this paper, the division of labor and scheduling among agents, and the superior performance of the LLMs.
4.1.2 Baselines
We employ proprietary and open-source LLMs in our experiments and set the temperature to 0.7 for proprietary models. For proprietary models, we adopt GPT-4o Achiam et al. (2023), and the O1-preview. For open-source LLMs, we adopt Qwen2.5-7B and Qwen2.5-14B Yang et al. (2024). We provide a detailed illustration of our designed prompts for different stages in Appendix A.
4.2 Experiment Results
4.2.1 Performance Comparison (RQ-1)
Topic-related Quality Evaluation: Table 2 illustrates the Topic-related Quality LAG achieved a recall of 67.58% and a precision of 70.33% with 20 items, indicating that it successfully retrieved a large proportion of relevant papers while maintaining a low rate of irrelevant ones. This performance is crucial for ensuring that the generated leaderboards are both comprehensive and accurate. The high precision and recall scores illustrate that LAG could help solve the paper coverage problem.
Fair Comparison: To ensure fair comparison, LAG extracted all experiment settings as part of the LAG-generated leaderboards. We provide a detailed case study of LAG-generated leaderboards with experiment settings in Appendix B.
Content Quality Evaluation: Table 2 presents the results of leaderboard quality generated by LAG and the baselines. LAG consistently achieved high scores across all evaluation metrics, particularly in terms of Coverage and Latest, indicating its ability to include a wide range of relevant and recent papers. For example, at a leaderboard length of 20 items, LAG achieved a Coverage score of 4.12 and a Latest score of 3.96, approaching human performance (4.72 and 4.68, respectively). While manual leaderboards scored slightly higher in content quality, LAG significantly reduced the time required for leaderboard generation, demonstrating its efficiency.
Figure 3: Impact of Iteration on LAG Performance.
Methods Leaderboard Length (items) Speed/s Content Quality Coverage Latest Structure Multiaspect LAG w/o Table Classification 5 $42.35$ $4.43$ $4.52$ $3.95$ $4.30$ LAG w/o Refinement $43.58$ $4.41$ $4.46$ $4.05$ $4.31$ LAG $49.64$ $4.52$ $4.70$ $4.32$ $4.51$ LAG w/o Table Classification 10 $81.37$ $4.31$ $4.36$ $4.01$ $4.33$ LAG w/o Refinement $80.52$ $4.24$ $4.17$ $3.88$ $4.10$ LAG $88.96$ $4.68$ $4.59$ $4.45$ $4.56$ LAG w/o Table Classification 15 $93.28$ $4.13$ $4.08$ $3.61$ $3.94$ LAG w/o Refinement $91.32$ $4.19$ $4.13$ $3.72$ $4.01$ LAG $105.61$ $4.47$ $4.16$ $4.32$ $4.28$ LAG w/o Table Classification 20 $105.31$ $3.92$ $3.88$ $3.51$ $3.77$ LAG w/o Refinement $99.35$ $3.85$ $3.71$ $3.62$ $3.72$ LAG $120.52$ $4.28$ $3.92$ $4.21$ $4.13$
Table 3: Ablation Study for LAG with different components removed. We use the GPT-4o as the backbone of LAG.
Iteration Evaluation: To ensure the high-quality of LAG-generated leaderboards, we iterated the process to evaluate the performance change during the whole iteration. Figure 3 presents the effect of different iteration counts on the performance of LAG. The results show that increasing the number of iterations from 1 to 5 provides a significant improvement in structure quality and Coverage quality scores. However, the latest score remains at a relatively high level, which is because in stage 1 of LAG, the old papers are filtered out.
Our experiments demonstrate that LAG is highly effective in generating high-quality, up-to-date leaderboards across various research topics. The framework’s ability to dynamically update leaderboards and extract detailed experiment settings ensures a fair comparison between state-of-the-art baselines. While LAG’s content quality scores are slightly lower than those of manually created leaderboards, its efficiency and scalability make it a valuable tool for researchers in rapidly evolving fields like AI and NLP.
4.2.2 Efficiency Analysis (RQ-2)
Construction Speed: LAG dramatically reduced the time required to generate leaderboards compared to manual methods. For instance, generating a 20-item leaderboard with LAG took approximately 120 seconds, while manual construction took over 18 minutes. This speed advantage makes LAG a practical tool for researchers who need up-to-date leaderboards in rapidly evolving fields. The high speed of LAG illustrates that LAG could help generate high-quality leaderboards timely.
4.2.3 Meta Evaluation (RQ-3)

### v2

@@ 3.3 Stage III Table Unpacking and Integration
Following the table extraction and classification phase, each table $\text{Table}_{i}$ is sent into the LLM to extract the core information. To build a useful and high-quality leaderboard, we define four types of scientific terms: Datasets, Metrics, Experiment Results, and Experiment Settings. We follow the definition of scientific entities proposed by SciIE  (Li et al., 2023). For datasets, we use LLMs to count the frequency in all filtered papers $P_{\text{filtered}}$ of each dataset under a certain research topic and retain the top-K (K=5) datasets with the highest frequency of occurrence in scientific papers, ensuring the generated leaderboard contains enough methods. For the rest of the three scientific terms, we utilize LLMs to extract from given $\text{Table}_{i}$ with a related table description $D_{i}$. After scientific term extraction, we recombine them into a quintuple, including the paper title as the unique identification ID. Each paper can produce one quintuple, and finally, we get a raw leaderboard with KMK*M quintuples from $M$ filtered papers and $K$ datasets. The raw leaderboard is then reranked on the basis of the experiment results.
3.4 Stage IV Leaderboard Generation and Evaluation
@@
Following the table unpacking and integration phase, we get KMK*M quintuples, and all quintuples are formatted into $K$ leaderboards. Each leaderboard is individually refined by a third-party LLM such as GPT-4 to enhance readability, eliminate redundancies, and ensure completeness. After we obtain $K$ leaderboards, the final stage involves a quality evaluation based on our pre-defined four criteria, which is shown in Appendix Table 9. Each leaderboard is assigned three scores based on “Coverage”, “Latest” and “Structure”. Since a research topic may contain several datasets, the “Multi-Aspect” is the average quality score used to evaluate the LLM-generated leaderboards for each dataset. The best leaderboard is chosen from $N$ candidates. LLMs critically examine the leaderboards in several aspects. The final output of League is $L_{\text{best}}=\text{Evaluate}({L_{\text{ca1}},L_{\text{ca2}},...,L_{\text{caN}}})$. The methodology outlined here, from paper collection to leaderboard evaluation, ensures that League effectively addresses the complexities of leaderboard generation in the AI domain using advanced LLMs. We provide Pseudo-code for easily understanding, which is shown in Appendix Algorithm 1.
Table 2: Results of leaderboard quality generated by League in the first iteration (The results of the first iteration could help reflect the effectiveness and efficiency of our framework). Items: The number of items in the leaderboard. For example, a 5-item leaderboard contains 5 baselines. Topic-related Quality: The precision and recall of each paper in relation to its relevance to the topic. Speed: The average time required to generate a single leaderboard. Content Quality: The evaluation results of the leaderboard content. The best results of Leagure are highlighted in bold, the fastest results are underlined.
[…]
4 Experiments
We designed experiments for League, with the aim of answering four questions: RQ-1: Can League address the paper coverage issue and generate fair leaderboards by incorporating the latest baselines? RQ-2: Can League reduce time consumption compared to human? RQ-3: Is the evaluation consistent between League and human experts? RQ-4: Is each proposed component of League useful?
4.1 Experimental Setup
4.1.1 Evaluation Metrics
We use two metrics to evaluate the quality (topic-related and leaderboard content) and the speed of leaderboard generation, respectively.
Table 3: Left: Table Classification result on the extracted tables using different LLMs. Right: Table NER result on the manually annotated table entities.
Methods Prompt P (%) R (%) F1 (%) GPT-4o 0-shot 80.96 82.49 80.03 1-shot 87.45 (+6.49) 86.44 (+3.95) 85.10 (+5.07) O1-preview 0-shot 81.16 80.23 78.55 1-shot 85.19 (+4.03) 83.62 (+3.39) 82.78 (+4.23)
Methods Prompt P (%) R (%) F1 (%) GPT-4o 0-shot 86.51 87.95 85.30 1-shot 89.17 (+2.66) 90.26 (+2.31) 88.76 (+3.46) O1-preview 0-shot 84.28 85.44 83.76 1-shot 88.55 (+4.27) 90.02 (+4.58) 89.82 (+6.06)
(1) Topic-related Quality: The aforementioned arXiv crawler employs regular expression matching in the abstract section to identify papers related to specific topics. While this method is efficient, it is relatively rudimentary and cannot guarantee that all retrieved papers meet our requirements. The quality of these papers not only directly affects the final leaderboard, but low-quality candidate papers can also significantly prolong the time required for construction. Therefore, it is essential to evaluate the quality of the retrieved articles. We evaluate the quality of content from the following two aspects. (i) Recall: It measures whether all items in the generated leaderboard are related to the given research topic. (ii) Precision: It identifies irrelevant items, ensuring that the items in the generated leaderboards are pertinent and directly support the given research topic.
(2) Leaderboard Content Quality: The evaluation metric of leaderboard Content Quality includes four aspects. Each aspect is judged by LLMs according to a 5-point, calibrated by human experts. The evaluation criteria are listed in Appendix Table 9. (i) Coverage: Each paper represented on the League-generated leaderboards encapsulates all aspects of the topic. (ii) Latest: Test whether all papers represented on the League-generated leaderboards are the latest. (iii) Structure: Evaluate the logical organization and determine whether the League leaderboards have missing items. (iv) Multiaspect: Average score of the previous three criteria for League-generated leaderboards.
(3) Leaderboard Construction Speed: Manually building a leaderboard is a time-consuming and laborious task. This process can be divided into the following main components: $T_{r}$ (search for papers on a specific topic), $T_{b}$ (browse all retrieved articles and develop several highly frequent data), $T_{f}$ (filter candidate papers based on the selected data), $T_{e}$ (read and extract information), and $T_{c}$ (the integration and construction time). The total time consumption is calculated as:
$T_{\text{manual}}=T_{r}+T_{b}+T_{f}+T_{e}+T_{c}.$ (1)
Given $L$ denotes the length of the leaderboard, $N_{\text{retrieved}}$ is the number of retrieved articles, $N_{\text{filtered}}$ is the number of articles retained, and $P$ is the proportion of valid articles with $P=\frac{N_{\text{filtered}}}{N_{\text{retrieved}}}$. We find that $T_{b}$ and $T_{f}$ are strongly correlated with leaderboard length $L$ and the Topic-related Quality:
Although $T_{r}$ is relatively fixed, $T_{e}$ and $T_{c}$ usually only have a positive correlation with $L$.
For League, we directly account for all the invocation time of the LLMs’ API calls. Compared to manual work, which often takes several days, League reduces the total time cost at the minute level. This is largely attributed to the task decomposition conducted in this paper, the division of labor and scheduling within the framework, and the superior performance of the LLMs.
4.1.2 Baselines
We employ proprietary and open-source LLMs in our experiments and set the sampling temperature to 0.7 for proprietary models. For proprietary models, we adopt GPT-4o (Achiam et al., 2023), and the O1-preview. For open-source LLMs, we adopt Qwen2.5-7B and Qwen2.5-14B (Yang et al., 2024). We provide a detailed illustration of our prompts for different stages in Appendix C.
4.2 Experiment Results
4.2.1 Performance Comparison (RQ-1)
Topic-related Quality Evaluation: Table 2 illustrates the Topic-related Quality League achieves a recall of 67.58% and a precision of 70.33% with 20 items, indicating that it successfully retrieves a large proportion of relevant papers while maintaining a low rate of irrelevant ones. The high precision and recall scores show that League can help solve the paper coverage problem. This performance is crucial to ensure that the generated leaderboards are both comprehensive and accurate.
Fair Comparison: To ensure a fair comparison, League extracts all experiment settings from crawled papers as part of the League-generated leaderboards. We provide a detailed case study of League-generated leaderboards with experiment settings in Appendix E.
Content Quality Evaluation: Table 2 and Table 5 in appendix present the results of the quality of the leaderboard generated by League. Items in Table 2 are crawled from arXiv while items in Table 5 are from the top-tier conferences (including ACL, EMNLP, NeurIPS, ICML, ICLR, etc.) respectively, which shows that League could guarantee the timelines and quality of the generated leaderboards. The papers crawled from arXiv could provide the latest research trend on the given topic, while papers crawled from top conferences could help researchers build leaderboards with higher quality, compared with the Topic-related Quality and Content Quality in Table 2 and Appendix Table 5. League consistently achieves high scores across all evaluation metrics, particularly in terms of Coverage and Latest, indicating its ability to include a wide range of relevant and recent papers. For example, at a leaderboard length of 20 items, League achieves a Coverage score of 4.12 and a Latest score of 3.96, approaching human performance (4.72 and 4.68, respectively). While manual leaderboards score slightly higher in Content Quality, League significantly reduces the time required for leaderboard generation from 1128s to 117.45s, demonstrating its efficiency. We also list a 5 item leaderboard generated by League, shown in Figure 3, which offers significant advantages over the outdated official HotpotQA leaderboard (Figure 1). Our method shows three advantages: 1) Latest, all papers are published after 2024; 2) High Quality, all papers are collected from top-tier conferences, including ACL, EMNLP, ICLR, and ICML; 3) More Information, all items on the leader board contain experiment details, including models and training strategies.
Table 4: Ablation study for League with components removed. We use the GPT-4o as the backbone.
[…]
Iteration Evaluation: To ensure the high-quality of League-generated leaderboards, we iterate the process to evaluate the performance change during the whole iteration. The left part of Appendix Figure 4 presents the effect of different iteration counts on the performance of League. The results show that increasing the number of iterations from 1 to 5 provides a significant improvement in Structure quality and Coverage quality scores. The Latest score remains at a relatively high level, which is because in stage 1 of the League, the old papers are filtered out. To sum up, our experiments demonstrate that League is highly effective in generating high-quality, up-to-date leaderboards across various research topics. The framework’s ability to dynamically update leaderboards and extract detailed experiment settings ensures a fair comparison between state-of-the-art baselines. While League’s Content Quality scores are slightly lower than those of manually created leaderboards, its efficiency and scalability make it a valuable tool for researchers in rapidly evolving fields like AI and NLP.
4.2.2 Efficiency Analysis (RQ-2)
Construction Speed: League dramatically reduces the time required to generate leaderboards compared to manual methods. For instance, generating a 20-item leaderboard with League takes approximately 2 minutes, while manual construction takes more than 18 minutes. This speed advantage makes League a practical tool for researchers who need up-to-date leaderboards in rapidly evolving fields. The high speed of League shows that it can generate high-quality leaderboards timely.
4.2.3 Meta Evaluation (RQ-3)
Figure 3: The example leaderboard generated by League. Comparing with the Leaderboard of Multi-hop QA method in the right part of Figure 1, our method could help summarize the experiment results from the top conference papers with experiment settings for fair comparison.
To verify the consistency between our proposed LLM evaluation strategy and human evaluation, we conduct a correlation evaluation involving human experts and our automated evaluation method. Human experts judge pairs of generated leaderboards to determine which one is superior. We compare the judgments made by our method against those made by human experts. Specifically, we provide experts with the same scoring criteria as used in our evaluation as a reference. Experts rank the 20 League-generated leaderboards and compare these rankings with those generated by LLM using the Pearson Correlation Coefficient to measure the consistency between human and LLM evaluations. The results of this meta-evaluation are presented in the right part of Appendix Figure 4. The table shows the Pearson Correlation Coefficient values, indicating a strong positive correlation between the quality scores provided by LLM and those given by human experts, with the O1-preview achieving the highest correlation at 0.76. These results suggest that our evaluation aligns well with human preferences, providing a reliable proxy for human evaluation.
4.2.4 Intermediate Evaluation
To validate the effectiveness of League’s intermediate stages, we conducted experiments focusing on table classification and table-level entity recognition, as summarized in Table 3. Table Classification. We divided tables into three categories: (i) main results, (ii) ablation studies, and (iii) others (e.g., dataset statistics or case studies). The results demonstrate that League is highly effective at isolating the correct “main results” tables required for leaderboard construction. Both GPT-4o and O1-preview provide strong performance (over 80% F1 scores respectively with 1-shot prompt). This suggests that even minimal supervision helps the model better distinguish between relevant and auxiliary tables, thereby reducing noise in the subsequent leaderboard generation stage. Table NER. Following Li et al. (2023), we cast entity extraction as a classification problem, requiring the model to identify four key scientific entities: methods, datasets, experimental settings, and metrics. On a manually annotated validation set, League shows robust performance across all categories, with notable improvements when enhanced prompting strategies are applied. These results confirm that the system can reliably extract structured information from experimental tables, which is critical for generating comparable high-quality leaderboards. Details of these tasks are shown in Appendix D.
4.2.5 Ablation Study (RQ-4)
To understand the contribution of each component in League, we conduct an ablation study by removing key components of League as follows: (1) League w/o Table Classification: We remove the table classification step, which leads to a slight decrease in Structure and Multiaspect scores, indicating that classifying tables is essential for maintaining a logical and well-organized leaderboard. (2) League w/o Refinement: We disable the Refinement step, which results in a minor drop in Coverage and Latest scores, suggesting that Refinement helps refine the leaderboard by ensuring that only the most relevant and recent papers are included. As shown in Table 4, the results confirm that each component plays a crucial role in achieving the generation of a high-quality leaderboard.
5 Conclusion
We introduced League, a novel framework leveraging LLMs to automatically generate the latest high-quality leaderboards based on given research topics. League addressed key challenges, including paper coverage, fair comparison, and timeliness, through a systematic approach that involves paper collection and splitting, table extraction and classification, table unpacking and integration, and leaderboard generation and evaluation. Experiments showed that League can automatically generate new high-quality leaderboards in a relatively short time and match human performance in terms of Topic-related Quality and Content Quality. This advancement offered a scalable and effective solution for synthesizing the latest leaderboards, providing a valuable tool for researchers in rapidly evolving fields such as artificial intelligence.
Ethics Statement
All authors affirm their adherence to the ICLR Code of Ethics. We have carefully considered the ethical implications of our research, particularly concerning the safe and responsible deployment of Large Language Model (LLM)s. Our work directly addresses the critical need to avoid harm by mitigating risks such as dangerous diagnostic medical recommendations, financial losses, and privacy breaches, which can arise from the unconstrained operation of LLMs. We believe our work contributes positively to human well-being by enhancing the safety and trustworthiness of advanced AI systems.
Reproducibility Statement
To ensure the reproducibility of our work, we have made significant efforts to document our methodology thoroughly. The full description and algorithm details of the League framework are described in Section 3 and algorithm 1. Our source code is provided in supplementary materials. We are committed to fostering open science and facilitating the replication of our results.
Appendix A Limitations
One limitation of League is its reliance on the quality of the retrieved papers. While our Topic-related Quality metrics are strong, there is still room for improvement in ensuring that all relevant papers are included. Future work could explore more sophisticated retrieval models to further enhance the coverage of the generated leaderboards. Another limitation is that a specific dataset may contain several evaluation metrics, and different papers may use different metrics to evaluate proposed models’ performance, bringing challenges for leaderboard generation and baseline comparison.
Appendix B The Use of Large Language Models
We employed LLMs for grammar checking and polishing the English expression throughout this manuscript. It is important to note that while our research focuses on leveraging LLMs for automatic leaderboard generation, the LLMs studied in this work are the subject of our research rather than tools for research ideation or scientific writing. All experimental design, analysis, and scientific conclusions were developed independently by the authors.
@@ Appendix B The Use of Large Language Models
1:Input: Scientific topic $T$, open-access platform arXiv AA
2:Output: Final refined and evaluated leaderboard $L$ # Stage 1: Paper Collection and Document Split
3: Crawl topic $T$ related NN publications $P_{\text{init}}=\{P_{1},...P_{N}\}\leftarrow\text{Retrieve}(T,A)$
4: Filter out topic-unrelated and old papers, $P_{\text{filtered}}=\{P_{1},...P_{M}\}\leftarrow\text{Retrieve}(P_{\text{init}},\text{date},\text{topic})$ # Stage 2: Table Extraction and Classification
5:for each Leaderboard iteration $i=1$ to ItersIters do
@@
13:   Recombine all quintuples, refine and rank the quintuples by performance scores.